Axolotl:

The following function represents the value of a car, in dollars, after x years: f(x) = 24,000(0.92)x What does 0.92 represent? The present value of the car The value of the car after x years The decrease in the value of the car, which is 92% The decrease in the value of the car, which is 8%

Axolotl:

I think it is The decrease in the value of the car, which is 92%

lifewmylilweeb:

how did you find this answer

Vocaloid:

You’re almost there - however, the base of the exponent for decay is actually 1 - r, so you’d need to set 1 - r = 0.92, and solve for r, the rate of decay

Vocaloid:

1 - r = 0.92 1 - 0.92 = r r = 0.08